OpenStudy (anonymous):

Please explain: xe^2+ye^2+6x-8y+20=0 represents a circle, find the center and radius.

OpenStudy (solomonzelman):

please correct the typos

OpenStudy (phi):

leave out the "e's" x^2+y^2+6x-8y+20=0 or \[ x^2+y^2+6x-8y+20=0 \]

OpenStudy (anonymous):

x squared+ y squared+6x-8y+20=0

OpenStudy (phi):

to find the center we write the equation as \[ (x^2+6x )+ (y^2-8y ) =-20 \] in other words, reordered so the x's are next to each other, and the same for the y's also, add -20 to both sides so that the "constant" is on the right-hand side.

OpenStudy (phi):

next, do you know how to "complete the square" ?

OpenStudy (anonymous):

no

OpenStudy (phi):

It is hard to explain here, but if you have time see https://www.khanacademy.org/math/algebra/quadratics/completing_the_square/v/solving-quadratic-equations-by-completing-the-square because the idea is complete the square for x^2 +6x and then do it again for y^2 -8y

OpenStudy (anonymous):

ok thank for your help i'll redo until i get it..thank again